### v2

# What Matters in Transformers? Not All Attention is Needed

Abstract
Scaling Transformer-based large language models (LLMs) has demonstrated promising performance across various tasks. However, this scaling also introduces redundant structures, posing challenges for real-world deployment. Despite some recognition of redundancy in LLMs, the variability of redundancy across different structures, such as MLP and Attention layers, is under-explored. In this work, we investigate the varying redundancy across different modules within Transformers, including Blocks, MLP, and Attention layers, using a similarity-based metric. This metric operates on the premise that redundant structures produce outputs highly similar to their inputs. Surprisingly, while attention layers are essential for transformers and distinguish them from other mainstream architectures, we found that a large proportion of attention layers exhibit excessively high similarity and can be safely pruned without degrading performance, leading to reduced memory and computation costs. Additionally, we further propose a method that jointly drops Attention and MLP layers, achieving improved performance and dropping ratios. Extensive experiments demonstrate the effectiveness of our methods, e.g., Llama-3-70B maintains comparable performance even after pruning half of the attention layers. Our findings provide valuable insights for future network architecture design. The code will be released at: https://github.com/Shwai-He/LLM-Drop.
What Matters in Transformers? Not All Attention is Needed
Shwai Hethanks: Equal contribution Guoheng Sun Zhenyu Shen Ang Lithanks: Corresponding author University of Maryland, College Park shwaihe@umd.edu, ghsun@umd.edu, zyshen@umd.edu, angliece@umd.edu
1 Introduction
Transformer-based large language models (LLMs) have significantly advanced AI research, achieving remarkable performance across various domains. Notable examples include ChatGPT OpenAI (2024) and Gemini Team (2024), which have demonstrated the potential of scaling LLMs to enhance capabilities in natural language understanding and generation. However, scaling these models introduces challenges, particularly the increase in redundant modules Frantar et al. (2023); Sun et al. (2023), which inflate deployment costs and resource demands. For instance, GPT-3, with its 175 billion parameters, requires approximately 350GB in FP16 format, far exceeding the 96GB memory capacity of the latest H100 GPUs, and making deployment on edge devices impractical. Recent studies Men et al. (2024); Gromov et al. (2024) have identified that many parameters in LLMs are redundant and can be pruned without compromising performance, highlighting the need to address redundancy within these models.
Although various previous works have been proposed to optimize LLM efficiency Frantar et al. (2023); Sun et al. (2023); He et al. (2024), these efforts often overlook the unique characteristics of transformer architectures. Specifically, transformer Vaswani et al. (2023) architectures are composed of multiple stacked blocks, each containing an MLP layer and an Attention layer, which exhibit different types and levels of redundancy. This motivates us to investigate the varying redundancy across these distinct modules.
Specifically, we propose a similarity-based metric to measure redundancy, operating on the premise that redundant modules produce outputs highly similar to their inputs. By computing the similarity between inputs and outputs, we can identify and drop redundant modules. Using this metric, we first empirically investigate Block Drop on the resulting performance. Even though some blocks exhibit relatively high similarity scores (e.g., more than 90% cosine similarity), Block Drop leads to substantial performance drops.
Given that a Transformer block comprises an MLP layer and an Attention layer, we examine the effects of dropping these layers separately. Our findings reveal that while dropping MLP layers negatively impacts performance, dropping Attention layers, i.e., the core of Transformer architectures which distinguish it from other mainstream architectures (e.g., RWKV Peng et al. (2023) and Mamba Gu and Dao (2024)), can be done extensively without degrading model performance. For example, dropping 50% of the Attention layers in Llama-2-70B Touvron et al. (2023) results in comparable performance to the original model, indicating a high level of redundancy in these layers.
To further enhance the effectiveness and efficiency of Layer Drop, we propose the Joint Layer Drop technique that jointly drops Attention and MLP layers, achieving higher dropping ratios and superior performance. Extensive experiments have demonstrated the effectiveness of our methods, e.g., Llama-3-70B maintains comparable performance even after pruning half of the attention layers. In short, our contributions are as follows:
  • We develop a similarity-based metric to measure the redundancy within LLMs and investigate the structured redundancy of various modules, including Blocks, MLP, and Attention layers.
  • We identify the surprising redundancy of Attention Layers and demonstrate the benefits of Attention Drop for accuracy, memory, and computational efficiency.
  • We propose additional techniques, e.g., Joint Layer Drop, to further enhance the effectiveness and efficiency of module dropping.
  • Our proposed Layer Drop method effectively identifies both important and redundant layers (e.g., excessive attention layers), which is crucial for existing model evaluation and provides valuable insights for future research on network architecture design.
2 Related Works
Large Language Models. Although Large Language Models (LLMs) have demonstrated promising performance across various tasks, their deployment costs still remain a significant challenge for practical usage. Transformer Vaswani et al. (2023) models consist of multiple blocks, which include attention layers and MLP layers. Attention layers compute the contextual information between input tokens with quadratic complexity concerning the input sequence length Li et al. (2020). KV-Cache Pope et al. (2022) mitigates the computational issue but results in excessive memory costs Zhang et al. (2023). MLP layers Liu et al. (2021); Mai et al. (2022) transform each token independently, using an up-projection followed by a down-projection, and account for most of the model’s parameters. Recent works have revealed that not all blocks or layers are important Men et al. (2024); Chen et al. (2024), which urges us to reflect on the architecture redundancy within LLMs and potentially more compact architectures.
Model Compression. LLMs can be compressed to promote their efficiency in memory and computation. Quantization Frantar et al. (2023); Lin et al. (2024) and Pruning Sun et al. (2023); Frantar and Alistarh (2023) are the most widely used techniques to compress LLMs. Specifically, quantization transforms the data type into low-bit but remains potentially redundant architecture and parameters. Pruning can be categorized into unstructured pruning and structured pruning, where the former does not apply to hardware while the latter degrades the performance significantly. Compared to pruning, our methods, i.e., Block Drop and Layer Drop, not only make the model architectures more compact but also maintain the performance. On the other hand, Block Drop and Layer Drop are orthogonal to quantization, and their integration with quantization significantly enhances the efficiency.
3 Methodology
In this section, we first discuss how to determine the redundancy in LLMs, and then propose corresponding algorithms to drop redundant modules.
Figure 1: Visualization of Layer Drop, where we visualize dropping either MLP or Attention Layers. Given the residual connection, we take LayerNorm together with the corresponding layers. The dropped layers with high similarity scores Refer to caption are blurred.
3.1 Similarity-based Drop
We utilize a similarity-based metric to determine the importance of modules within LLMs by computing the similarity between the input and output of the corresponding modules. The underlying motivation is that redundant modules produce outputs similar to the inputs, so skipping such modules does not significantly degrade performance. Conversely, important modules are expected to transform the inputs substantially and cannot be skipped. Therefore, we use the cosine similarity between the hidden states of the input and output to represent redundancy. Given the input $\boldsymbol{X}$ and output $\boldsymbol{Y}$, the importance score can be measured as follows:
$\boldsymbol{S}=1-Cosine(\boldsymbol{X},\boldsymbol{Y}),$ (1)
where Cosine denotes the cosine similarity. Higher cosine similarity would lead to lower importance scores, and the modules with the lowest scores would be dropped.
Figure 2: Visualization of Block Drop, where we use the annotation Refer to caption to highlight the layers with high similarity scores between the corresponding inputs and outputs. The dropped blocks are blurred.
3.2 Block Drop
Transformer models are always stacked by multiple blocks, where each block shares the same architecture and can be seen as a subnetwork. Therefore, we first consider dropping unimportant blocks.
As shown in Figure 2, the transformer architecture operates sequentially, where the output of each block is fed into the next block. Based on this workflow, we directly compute the similarity between the inputs and corresponding outputs. Specifically, for the $l$-th block, the importance score is measured as follows:
$\boldsymbol{S}^{l}_{{B}}=1-Cosine(\boldsymbol{X}^{l}_{B},\boldsymbol{Y}^{l}_{B}),$ (2)
where $\boldsymbol{X}^{l}_{B}$ and $\boldsymbol{Y}^{l}_{B}$ denote the input and output of the $l$-th block, respectively. Since the similarity scores are computed locally, we can offload irrelevant modules to save memory. Specifically, we iteratively compute the importance score of each block from shallow to deep, and then drop the blocks with the lowest scores.
3.3 Layer Drop.
A transformer block consists of an Attention layer and an MLP layer, each performing different functions on the input sequence, i.e., contextual information flow between tokens and token transformation. Considering the different functionalities of Attention and MLP, we discuss them separately.
MLP Drop. As illustrated in Figure 1, an MLP layer follows LayerNorm and its outputs are a part of the overall output which involves residual connections. Given the input $\boldsymbol{X}^{l}_{M}$ of the LayerNorm before MLP at the $l$-th Block, the output $\boldsymbol{Y}^{l}_{{M}}$ can be formulated as:
$\boldsymbol{Y}^{l}_{{M}}=\boldsymbol{X}^{l}_{{M}}+\text{MLP}(\text{LayerNorm}(\boldsymbol{X}^{l}_{{M}})),$ (3)
where $\boldsymbol{Y}^{l}_{{M}}$ involves two parts, and the output of the MLP constitutes only a portion of it. Therefore, directly utilizing the similarity metric on the inputs and outputs of the MLP leads to inaccurate measurements. To capture the overall inputs and outputs accurately, we consider the MLP layers together with the subsequent LayerNorm as a single unit for dropping. This approach ensures a more accurate measurement. The importance metrics are then as follows:
$\boldsymbol{S}^{l}_{{M}}=1-Cosine(\boldsymbol{X}^{l}_{M},\boldsymbol{Y}^{l}_{M}).$ (4)
Note that we consider MLP layers together with their corresponding LayerNorm layers. Therefore, MLP Drop removes both the insignificant MLP layers and the associated LayerNorm layers.
Attention Drop. Similar to MLP Drop, Attention Drop also considers the residual connection, as the outputs of attention layers contribute only partially to the overall outputs:
$\boldsymbol{Y}^{l}_{{A}}=\boldsymbol{X}^{l}_{{A}}+\text{Attention}(\text{LayerNorm}(\boldsymbol{X}^{l}_{{A}})),$ (5)
where $\boldsymbol{X}^{l}_{{A}}$ is the inputs of the corresponding LayerNorm layers and $\boldsymbol{Y}^{l}_{{A}}$ overall outputs that involves residual connections. Like MLP Drop that takes attention layers and associated LayerNorm layers together, the importance score for Attention Drop is:
$\boldsymbol{S}^{l}_{{A}}=1-Cosine(\boldsymbol{X}^{l}_{A},\boldsymbol{Y}^{l}_{A}).$ (6)
Similar Block Drop, the computation of importance scores of Layer Drop is conducted layer-wisely from shallow layers to deep layers. All Drop methods are conducted with a calibration dataset $\mathcal{D}$ (e.g., C4 Raffel et al. (2020)) in a one-shot manner that computes the importance scores once and then drops redundant layers, instead of using resource-consuming multi-turn iterations. The effectiveness of One-Shot Drop has been verified in Appendix B. On the other hand, other than solely dropping Attention or MLP layers individually, we propose Joint Layer Drop that considers the combined importance scores across these layers to optimize the dropping process. The algorithm for Joint Layer Drop is presented in Algorithm 1.
Initialize: $\boldsymbol{S}\leftarrow\emptyset$.
for $l=1$ to $L$ do
 Compute $\boldsymbol{S}_{M}^{l}$ and $\boldsymbol{S}_{A}^{l}$ via Eq. 4 and 6.
$\boldsymbol{S}\leftarrow\boldsymbol{S}\cup\{\boldsymbol{S}_{M}^{l},\boldsymbol{S}_{A}^{l}\}$.
end for
Set the threshold $\tau$ to the k-th element of $\boldsymbol{S}$.
Drop the layers with scores lower than $\tau$.
Algorithm 1 Joint Layer Drop
After removing the insignificant experts, the pruned model can be easily loaded using existing packages (e.g., Huggingface Transformers Wolf et al. (2020)) with just a change of the model configuration. More implementation details are listed in the Appendix A.
(a) Block Drop.
(b) MLP Drop.
(c) Attention Drop.
Figure 3: Visualization of Dropping order for Block Drop and Layer Drop. We visualize of the remaining layers and blocks under different dropped numbers, where yellow areas represent the retained portions and red areas indicate the dropped layers/blocks.
Table 1: Experimental results of dropping different modules on Llama-2-13B and Mistral-7B, where Block, MLP and Attn are corresponding module. We drop a fixed number of modules (i.e., 4 and 8), denoted by the postfix "-$n$" ($n$ refers to the number of dropped modules). “Memory” refers to the memory cost associated with deploying models. Rows with averaged performance lower than 95% of the original performance are grayed.
Llama-2-13B
Method SpeedUp Memory ARC-C BoolQ HellaSwag MMLU OBQA PIQA RTE WinoGrande Avg.
Baseline 1.00×1.$0$ 24.4GB 59.9 80.7 82.2 55.1 45.6 80.5 65.0 77.0 68.2
Block-4 1.13×1.13$\times$1.13 × 22.0GB 54.8 73.3 80.6 54.8 45.8 79.1 60.3 77.5 65.8
Block-8 1.23×1.23$\times$1.23 × 19.6GB 48.0 56.8 75.3 53.8 41.2 75.3 59.9 75.6 60.7
MLP-4 1.05×1.05$\times$1.05 × 22.8GB 54.9 76.1 80.4 54.8 45.4 79.5 66.4 77.3 66.9
MLP-8 1.09×1.09$\times$1.09 × 21.2GB 49.2 63.4 75.6 54.5 42.2 76.0 59.2 75.1 61.9
Attn-4 1.05×1.05$\times$1.05 × 23.6GB 58.8 80.4 82.0 54.7 46.2 80.5 67.9 77.2 68.5
Attn-8 $1\times1$ 22.8GB 58.2 80.5 82.2 54.5 47.0 80.5 64.3 77.4 68.1
Attn-16 1.23×1.23$\times$1.23 × 21.3GB 56.4 79.2 81.9 48.2 47.4 79.5 59.9 76.2 66.1
Attn-20 1.30×1.30$\times$1.30 × 20.5GB 53.8 76.9 78.6 51.5 44.4 77.6 59.2 77.1 64.9
Mistral-7B
Method SpeedUp Memory ARC-C BoolQ HellaSwag MMLU OBQA PIQA RTE WinoGrande Avg.
Baseline 1.00×1.$0$ 14.0GB 61.5 83.7 83.2 62.5 43.8 82.0 66.8 78.5 70.3
Block-4 1.12×1.12$\times$1.12 × 12.3GB 53.1 80.4 77.5 61.6 40.0 77.6 70.0 76.6 67.1
Block-8 1.29×1.29$\times$1.29 × 10.6GB 40.0 71.6 63.9 60.0 30.6 69.3 63.9 69.7 58.6
MLP-4 1.05×1.05$\times$1.05 × 12.7GB 53.2 80.3 77.7 61.7 40.0 77.6 67.5 77.3 66.9
MLP-8 1.09×1.09$\times$1.09 × 11.3GB 36.7 71.8 33.6 53.3 30.6 68.0 66.8 66.6 53.4
Attn-4 1.09×1.09$\times$1.09 × 13.6GB 61.0 83.5 82.9 62.5 44.6 82.0 64.6 78.0 69.9
Attn-8 1.17×1.17$\times$1.17 × 13.2GB 60.2 82.7 82.3 62.2 44.2 81.3 66.8 78.8 69.8
Attn-12 1.24×1.24$\times$1.24 × 12.8GB 57.2 76.8 80.2 59.4 41.8 79.1 66.1 77.7 67.3
Figure 4: Visualization of Importance Scores in Checkpoints during the Pre-training Process of MAP-Neo-7B, where lighter areas represent low importance scores (i.e., high similarity scores). We present both (1) Early Stage (the first 12 released checkpoints) and (2) Entire Process (checkpoints for every 500B trained tokens). We independently visualize the importance scores at the module index 0, since they are significantly higher.
Table 2: Block Drop and Layer Drop on Larger Models, where we drop a series of numbers (from 4 to 48) of modules on Llama-2-70B.
Llama-2-70B
Method SpeedUp Memory ARC-C BoolQ HellaSwag MMLU OBQA PIQA RTE WinoGrande Avg.
Baseline 1.00×1.$0$ 128.7GB 67.4 83.8 87.1 68.5 48.6 82.5 69.3 83.7 73.9
Block-4 1.10×1.10$\times$1.10 × 122.4GB 65.7 83.9 86.2 68.9 47.0 82.3 68.2 84.0 73.3
Block-8 1.21×1.21$\times$1.21 × 116.1GB 61.9 83.8 85.0 68.9 47.8 81.8 65.7 82.3 72.2
Block-16 1.32×1.32$\times$1.32 × 103.3GB 56.3 83.9 81.2 69.2 46.2 79.0 61.4 80.7 69.7
Block-32 1.56×1.56$\times$1.56 × 77.7GB 41.1 63.0 59.0 65.6 33.0 64.7 66.4 67.6 57.6
MLP-4 1.05×1.05$\times$1.05 × 123.5GB 65.4 84.0 86.1 68.7 46.6 82.9 68.2 83.4 73.2
MLP-8 1.14×1.14$\times$1.14 × 118.3GB 64.4 83.9 84.9 68.7 47.6 81.7 66.8 82.2 72.5
MLP-16 1.24×1.24$\times$1.24 × 107.8GB 57.5 53.6 81.6 69.1 46.0 79.2 58.8 81.7 65.9
MLP-32 1.41×1.41$\times$1.41 × 86.78GB 40.6 61.9 64.2 59.8 29.8 64.2 52.7 72.7 55.7
Attn-4 1.04×1.04$\times$1.04 × 127.7GB 67.2 84.0 87.0 68.6 48.8 82.5 69.3 83.3 73.8
Attn-8 1.08×1.08$\times$1.08 × 126.5GB 67.3 83.8 86.9 68.5 48.4 82.9 69.0 82.6 73.7
Attn-16 1.12×1.12$\times$1.12 × 124.3GB 67.8 83.9 87.2 68.5 49.0 83.0 68.2 82.8 73.8
Attn-32 1.16×1.16$\times$1.16 × 119.7GB 67.2 84.8 87.2 68.4 49.6 81.8 67.5 83.5 73.8
Attn-40 1.21×1.21$\times$1.21 × 117.5GB 63.7 82.8 84.4 66.2 46.8 80.1 66.8 81.3 71.5
Attn-48 1.23×1.23$\times$1.23 × 115.3GB 58.5 73.7 80.6 56.8 45.0 79.8 59.6 81.0 66.9
4 Experiments
4.1 Main Results
Comparison of Dropping Different Target Modules. In Table 1, we compare dropping different modules (i.e., Block, MLP, and Attention) with fixed numbers (e.g., 4 and 8). Since Blocks are the fundamental components of transformers, we first investigate Block Drop. Block Drop significantly degrades the performance, e.g., decreasing by 2.4% and 7.5% on the averaged performance when dropping 8 Blocks on Llama-2-13B, respectively. Given the large performance gap, Block drop is sufficient to maintain the performance.
We next delve into removing fine-grained modules in Blocks, i.e., MLP and Attention Layers. Similar to Block Drop, dropping MLP layers also leads to a substantial performance decline, e.g., 6.3% on Llama-2-13B and 6.9% on Mistral-7B when dropping 8 layers, which is unacceptable for practical usage.
Surprisingly, dropping attention layers demonstrates high effectiveness, as models with 8 dropped layers still maintain most of their performance, e.g., both Llama-3-13B and Mistral-7B maintain more than 99% of the original performance after dropping 8 attention layers. Although attention layers are the hallmarks of transformers, dropping such layers does not degrade the performance.
Efficiency Promotion of Module Dropping. We also compare the practical speedup memory usage of dropping different modules. We run models with various input and generated sequence lengths (from 512 to 4096) on a single Nvidia A100 GPU and report the average cost. While dropping blocks that contain both attention layers and MLP layers offer the best efficiency, the significant performance drop limits practical application. Compared to MLP layers, attention layers involve fewer memory costs but have more complex operations, resulting in greater contributions to speedup after dropping (1.17×1.17$\times$1.17 × v.s. 1.09×1.09$\times$1.09 × when dropping 8 layers on Mistral-7B).
(a) Evaluation of Block Drop.
(b) Evaluation of MLP Drop.
(c) Evaluation of Attention Drop.
Figure 5: Curves of Performance with respect to Dropping Ratios. The solid lines represent the impact of dropping the $n$ modules with the lowest importance scores in Mistral-7B and Llama-2-13B, and the dotted lines represent the performances of the baseline and random guessing.
Attention Drop Maintains the Performance. To obtain a universal understanding of module dropping, we further conduct experiments on different a series of dropping ratios (from 0% to 100%), and the results are shown in Figure 5. Transformer models exhibit sensitivity to both Block Drop and Layer Drop, where even dropping a small number of blocks or MLP layers results in a significant performance decline. However, Attention Drop initially shows a stable trend that maintains comparable performance when dropping ratios are below 50% but subsequently decreases due to excessive layer removal.
Attention Drop Significantly Reduces the KV Cache. Given the auto-regressive nature of attention, where the output is generated token by token, the key-value (kv) cache is used to store intermediate representations of input sequences to speed up the inference process by avoiding redundant computations. Despite its role in speeding up inference, the kv cache results in significant extra memory usage Zhang et al. (2023); Singhania et al. (2024). For instance, an input with a batch size of 128 and a sequence length of 2048 for Llama-2-13B would result in about 70GB of kv cache. Our proposed Attention Drop method can safely remove 20 out of 40 attention layers, effectively reducing the kv cache by half and significantly lowering the memory requirements.
4.2 Detailed Analysis
Consistent Redundancy of Attention Layers Throughout Training Progression. To investigate how high similarity patterns are achieved, we revisit the historical checkpoints to track the dynamic changing of similarity.
Specifically, we evaluate checkpoints released by MAP-Neo-7B Zhang et al. (2024), since it released continuous checkpoints during training stages. Figure 4 presents the importance scores of Blocks and Layers at different training stages, where Attention layers demonstrate excessively low importance scores at all training stages. However, for MLP layers and Blocks, the importance scores gradually increase as training progresses, indicating that they cannot be dropped as readily as Attention layers.
Table 3: Experimental results on Llama-3, where Llama-3-8B and Llama-3-70B are included.
 Method HellaSwag  MMLU  OBQA  WinoGrande   Avg.
Llama-3-8B
Baseline 82.2 65.5 45.0 77.7 67.6
Attn-4 81.6 65.1 44.8 78.2 67.4
Attn-8 81.1 65.1 45.0 78.4 67.4
Attn-12 79.4 63.9 42.2 77.8 65.8
Attn-16 71.2 38.2 39.4 72.8 55.4
Attn-20 42.2 23.0 30.6 58.7 38.6
Llama-3-70B
Baseline 88.0 78.7 48.4 85.4 75.1
Attn-4 87.9 78.7 49.0 85.2 75.2
Attn-8 87.8 78.5 48.8 85.2 75.1
Attn-16 87.8 78.7 48.6 84.9 75.0
Attn-32 87.9 78.6 48.8 85.3 75.2
Attn-40 85.2 77.1 48.0 82.8 73.3
Attn-48 81.2 73.9 47.4 81.3 71.0
Deeper Modules Display Higher Redundancy. To further analyze the dropped modules, we visualize the dropped layers or blocks. Figure 3 visualizes the remaining and dropped layers/blocks as the number of dropped modules increases. Llama-2-13B and Mistral-7B exhibit similar patterns in Layer Drop and Block Drop: initially, both models tend to drop the deeper layers, followed by the shallower ones. These findings are consistent with Xu et al. Men et al. (2024), which suggests that deeper layers tend to be more redundant. Larger models (e.g., Llama-2-70B) also showcase a similar trend, which is shown in Appendix B.
(a) Llama-2-13B
(b) Mistral-7B
Figure 6: Accuracy Curves of Dropping Different Target Modules, where we consider dropping single types of modules and Joint dropping (Attn + MLP). In the Joint’s line, $\bigstar$ represents the step where the MLP is dropped, while the $\blacksquare$ represents the step where the Attention is dropped.
Block Drop and Layer Drop are Data Robust. In Figure 7 and Figure 11, we note that feature similarity remains relatively stable across different Attention layers as the sample size increases (i.e., from 4 to 1024), indicating that Block Drop and Layer Drop maintain consistency regardless of sample quantity. This confirms that using $256$ samples suffice for computing similarity, which is adopted for all our experiments. Similarly, Figure 8 and Figure 12 show that varying the calibration datasets, from the pretraining dataset (e.g., C4 Raffel et al. (2020)) to instruction tuning datasets (e.g., CodeAlpaca-20k [^1] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k, MathInstruct Yue et al. (2024) and LIMA Zhou et al. (2024)), does not significantly alter feature similarity. Despite some differences, the modules dropped remain the same. This demonstrates the resilience of Block Drop and Layer Drop to variations in calibration datasets.
Figure 7: The Impact of Sample Quantity on the Importance Scores of Attention Layers.
Figure 8: The Impact of Dataset Type on the Importance Scores of Attention Layers.
Larger Models are more Robust to Dropping. To verify the consistency of our findings on larger models, we take Llama-2-70B into consideration, since it also comes from the Llama family and larger model size. Specifically, we drop the modules with different dropping ratios ranging from 5% to 60% on Table 2.
Similar to the findings in smaller models, Llama-2-70B also showcases sensitivity to Block Drop and Layer Drop, where dropping only 20% of blocks or MLP layers leads to a significant performance drop. However, it is noted that larger models are more robust than smaller models: Llama-2-70B dropped with 8 MLP layers only encounters 0.6% performance drop on average, compared to 3.9% of Llama-2-13B when dropping 4 MLP layers. Table 3 also reveals a similar trend on Llama-3.
On the other hand, Attention Drop performs well on Llama-2-70B. Specifically, when dropping 32 out of 80 attention layers, the performance of Llama-2-70B decreases by only 0.1% on average. Conversely, on Llama-2-13B, dropping the same proportion of attention layers leads to a noticeable performance drop.
Joint Layer Drop Enhances the Performance. In this part, we assemble two strategies of Layer Drop, i.e., Attention Drop and MLP Drop. Specifically, we first obtain the importance scores ($\boldsymbol{S}_{{A}}$ and $\boldsymbol{S}_{{M}}$) for Attention layers and MLP Layers, respectively. Then, we concatenate the importance scores ($\boldsymbol{S}=[\boldsymbol{S}_{{A}},\boldsymbol{S}_{{M}}]$), and drop the layers with the lowest scores.
As shown in Figure 6, the performance of Joint Drop is noticeably higher than either Attention Drop or MLP Drop alone. Joint Drop first only drops Attention Layers until the number of dropped layers exceeds 14 for Mistral-7B and 18 for Llama-2-13B. This is why the curves for Attention Drop and Joint Drop overlap in the left part of the graph. This process demonstrates that Attention layers are more redundant than MLP layers, which is consistent with the differing performance observed between MLP Drop and Attention Drop. Notably, Joint Drop improves the performance of Attention Drop, achieving higher dropping ratios and superior performance.
Insights for Future Network Architecture Design. Despite the success of scaling up large language models, our work provides valuable insights for scaling down models to achieve more efficient architectures. On the one hand, given the high redundancy of attention layers, future works can reduce the number of attention layers (e.g. dropping deeper attention layers), rather than keeping it consistent with that of MLP layers. On the other hand, unlike MLP layers, attention layers demonstrate consistent redundancy as training progresses, which may be a potential bottleneck in training large language models. Replacing attention layers or developing new training techniques may further enhance the capacity of language models.
5 Conclusion
In this work, we systematically revisit transformer architectures through the lens of module dropping, exploring three different structures: Blocks, MLP layers, and Attention layers. Our findings reveal that attention layers exhibit surprisingly high redundancy and can be removed in large proportions without compromising performance. Additionally, we further propose Joint Layer Drop, which further promotes the dropping ratios and performance. Our study empirically demonstrates the potential to make transformer models more compact and efficient, providing valuable insights for future network design within the NLP community.
6 Limitations
Despite the progress we have made, there are still limitations in our work. While our discussed dropping techniques demonstrate improved efficiency on the models we examined, a potential future work is a further exploration to verify their application across a broader range of models, e.g., vision transformers and vision-language models. Additionally, our work mainly focuses on post-training dropping and does not involve retraining, which has the potential to recover or even enhance performance further. Our work provides valuable insights for network architecture design, and we believe that incorporating further training will contribute to superior performance in more compact architectures.
Appendix A Implementation Details
Models.
We utilize Llama-2 Touvron et al. (2023) and Mistral Jiang et al. (2023) as the default models, given their competitive performance and wide usage. We also evaluated the completely open-source model MAP-Neo Zhang et al. (2024) to explore the redundancy variations in the modules of the entire pre-training phase. Additionally, we experimented with the newly released Llama-3 to verify the effectiveness of model dropping on the latest models.
Datasets.
For the calibration dataset, we used the validation set of C4 dataset Raffel et al. (2019), with 256 samples and an input sequence length of 2,048, following the setup in Sun et al. (2023). To evaluate the performance of the model, we report the results of the following tasks: BoolQ Clark et al. (2019), OBQA Mihaylov et al. (2018), PIQA Bisk et al. (2019), RTE Wang et al. (2019), ARC-C Clark et al. (2018), HellaSwagZellers et al. (2019), MMLU Hendrycks et al. (2021), WinoGrande ai2 (2019) and GSM8K Cobbe et al. (2021). Please refer to Table 4 for detailed information. The evaluation code is based on EleutherAI LM Evaluation Harness Gao et al. (2023).
Table 4: Parameter settings for each task. “Norm” refers to the normalization performed with respect to the length of the input.
 Task  Number of few-shot  Metric
BoolQ 0 Accuracy
RTE 0 Accuracy
OBQA 0 Accuracy (Norm)
PIQA 0 Accuracy (Norm)
MMLU 5 Accuracy
WinoGrande 5 Accuracy
GSM8K 5 Exact Match
HellaSwag 10 Accuracy (Norm)
ARC-C 25 Accuracy (Norm)
Appendix B Additional Experimental Results
Figure 9: Accuracy Curves on GSM8k, where Mistral-7B-Base represents mistralai/Mistral-7B-v0.1, Llama-2-7B-Base represents meta-llama/Llama-2-7b-hf, Mistral-7B-Math represents meta-math/MetaMath-Mistral-7B, and Llama-2-7B-Math represents meta-math/MetaMath-7B-V1.0.
Performance on Tasks Requiring Specialized Knowledge.
To evaluate Attention Drop on more complex technical tasks, we evaluated Llama-2-7B and and Mistral-7B, and two corresponding instruction fine-tuned models, meta-math/MetaMath-7B-V1.0 and meta-math/MetaMath-Mistral-7B Yu et al. (2023). The results in Figure 9 indicate that, except for Llama-2-7B-Math, which is meta-math/MetaMath-7B-V1.0, all the models do not experience significant performance degradation when dropping fewer than 8 Attention layers. We speculate that this is because Llama-2-7B-Math is initialized with Llama-2-7B and undergoes instruction fine-tuning to improve its mathematical ability. Llama-2-7B-Base exhibits poor performance in mathematics, and the ability obtained solely through fine-tuning appears to be superficial. Therefore, when dropping Attention layers, Llama-2-7B-Math’s ability rapidly deteriorates.
We select the first sample from the test set of GSM8K [^2] https://huggingface.co/datasets/openai/gsm8k as the input and display the model’s raw output in Table 6. Even when dropping 10 attention layers, the model could still compute the correct answer of this question, but it failed to adhere to the correct output format. However, when dropping 12 attention layers, the model is no longer able to produce the correct answer of this question.
Table 5: Integration of Module Dropping and Quantization. “w/Quant” denotes quantized models.
 Method  ARC-C  HellaSwag  OBQA  WinoGrande  Avg.
Llama-2-13B
Baseline 59.9 82.2 45.6 77.0 66.2
w/Quant 59.5 81.7 45.8 77.1 66.0
Attn-4 58.8 82.0 46.2 77.2 66.1
w/Quant 58.0 81.7 46.0 76.2 65.5
Attn-8 58.2 82.2 47.0 77.4 66.2
w/Quant 57.7 81.9 47.0 77.0 65.9
Mistral-7B
Baseline 61.5 83.2 43.8 78.5 66.8
w/Quant 61.2 82.5 42.8 78.0 66.1
Attn-4 61.0 82.9 44.6 78.0 66.6
w/Quant 61.0 82.8 43.6 77.6 66.3
Attn-8 60.2 82.3 44.2 78.8 66.4
w/Quant 60.1 82.0 43.8 77.5 65.9
Table 6: Output Cases of Model with Dropping Some Attention Layers, where correct answers are marked in green, while incorrect answers are marked in red. Note that “####” is used to extract the correct answer.
Question Janet’s ducks lay 16 eggs per day. She eats three for breakfast every morning and bakes muffins for her friends every day with four. She sells the remainder at the farmers’ market daily for \$2 per fresh duck egg. How much in dollars does she make every day at the farmers’ market?
[…]
Method Output
[…]
Simple One-Shot Drop is Effective.
Iterative and One-Shot are the most common settings for model compression, where the former prunes step by step while one-shot prune the model once. Iterative pruning may causes more accurate measurements but tend to be more time-consuming. In Figure 10, we empirically compare Iterative Dropping and One-Shot Dropping. For Iterative Dropping, we consider dropping one by one iteratively.
However, based on Figure 10, Iterative Dropping achieves performance that is merely comparable to One-Shot Dropping, without offering any significant enhancement. Given its simplicity and efficiency, One-Shot Dropping emerges as the superior choice.
Figure 10: Ablation study on dropping strategies, i.e., Iterative and One-Shot, where One-Shot Dropping achieves comparable performance with Iterative Dropping.
Module Dropping is Orthogonal to Quantization.
Given that quantization simplifies data types and enhances efficiency in memory usage and inference speed, we integrate module dropping with quantization to verify whether the quantized models can maintain the performance achieved by Attention Drop. Specifically, we use the mainstream AWQ algorithm Lin et al. (2024) for 4-bit quantization, following its default settings, which involve using 128 samples from the Pile dataset Gao et al. (2020) as the calibration dataset.
As shown in Table 5, the integration of quantization still maintains the performance of Attention Drop, i.e., only less than 1% difference in average performance.
Robustness to Calibration Datasets.
Based on Figure 11 and 12, similar to Attention Drop, the importance scores for MLP Drop and Block Drop also demonstrate robustness to calibration datasets, regardless of changes in data type or sample size. Although minor differences exist in the scores, the resulting dropped modules are nearly identical.
Dropping Order on Larger Models.
We present the dropping order of Block Drop and Layer Drop for the 70B Llama models in Figure 13. Similar to smaller models, larger models also tend to drop deeper layers first. While the dropping order of Blocks differs between Llama-2-70B and Llama-3-70B, we believe this is attributed to different training techniques, e.g., different number of training tokens.
(a) Block
(b) MLP
Figure 11: The Impact of Sample Quantity on the Importance Scores of Block and MLP.
(a) Block
(b) MLP
Figure 12: The Impact of Dataset Type on the Importance Scores of Block and MLP.
(a) Block
(b) MLP
(c) Attention
Figure 13: Visualization of Dropping order for Block Drop and Layer Drop on Larger Models, i.e., Llama-2-70B and Llama-3-70B.